OpenStudy (anonymous):

The Pretty Pizza Company wants to open a new pizza place either on the north side of town or the south side of town. They want to open the pizza place where the ratio of pizza places per person is currently less than 0.008. The population of 15 blocks of the north side of town is 220 people per block, but there are already 26 pizza places in the area. The population of 22 blocks of the south area is 175 people per block, and there are 33 pizza places in the area. Where, if anywhere, should the Pretty Pizza Company open a new pizza place? South side North side Both sides niether

OpenStudy (anonymous):

On the north side there are 220 people per block and there are 15 blocks so you do 220*15= 3,300. But then there are already 26 pizza placed so you take 26/3,300 to get the ratio of .0078 pizza places per one person. And on the south side there are 175 people for 22 block, so it's 175*22= 3,850. Like before you take the amount of pizza places and divide it by the amount of people, so it's 33/3,850=.0085 pizza places per one person. Now, Pretty Pizza Place wanted to put a new restaurant where the ratio of pizza place per one person was less than .008 so they should go to the north side.